1. Calculate the Price Elasticity of Demand for FIRM 1, ηd, between the following two points: (use the midpoint formula)

Price of Good Quantity of Good
24 1,900
41 1,520
Answer #1

Answer
Elasticity of demand=(change in quantity/average quantity)/(change in price/average price)
Change in quantity=1520-1900=-380
average quantity=(1520+1900)/2=1710
change in price=41-24=17
average price=(41+24)/2=32.5
Elasticity of demand=(-380/1710)/(17/32.5)
=-0.42
the elasticity of demand is -0.42 in absolute terms
It is below 1 so the demand is inelastic.